A ball with a mass of 1 gramme and a charge of 10^-8 C moves from point A whose potential is 600 V to point B whose potential is zero. What was the velocity of the ball at point A if at point B it is 20 cm/s?



1
Expert's answer
2022-01-31T07:34:15-0500

Answer

Potential difference

"V_{AB}=600-0=600volts."

Charge Q="10^{-8}C"

Speed of B

"v_B=0.2m\/s"

Using energy conservation

"QV_{AB}=0.5m(v_A^2-v_B^2)"

Putting all values

"v_B^2=\frac{10^{-8}\times600}{0.5\times0.001}+0.04\\=0.07\\v=0.26m/sec"